OpenStudy (anonymous):

find the exact value by using half - angle identity sing pi/8 Help plzz

OpenStudy (jdoe0001):

\(\bf sin\left(\cfrac{\pi}{8}\right)\\ \quad \\ \quad \\ 2\times \cfrac{1}{8}\implies \cfrac{1}{4}\qquad thus\\ \quad \\ sin\left(\cfrac{\pi}{8}\right)\implies sin\left(\cfrac{\frac{\pi}{4}}{2}\right)\) use that in your half-angle trig identity for sine

OpenStudy (anonymous):

\[\sin \pi/4/2= -+\sqrt{1-cosx}/2\]

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

i cant finish it

OpenStudy (jdoe0001):

.. why not?

OpenStudy (anonymous):

what i have to used for cos x

OpenStudy (jdoe0001):

\(\bf sin\left(\cfrac{\frac{\pi}{4}}{2}\right)=\sqrt{\cfrac{1-cos\left(\frac{\pi}{4}\right)}{2}}\) that angle will be in the Unit Circle

OpenStudy (anonymous):

\[-+\sqrt{1-\sqrt{2}/2 }/2\]

OpenStudy (jdoe0001):

ahemm so... \(\bf sin\left(\cfrac{\frac{\pi}{4}}{2}\right)=\sqrt{\cfrac{1-\frac{\sqrt{2}}{2}}{2}}\implies \sqrt{\cfrac{\frac{2-\sqrt{2}}{2}}{2}}\implies \sqrt{\cfrac{2-\sqrt{2}}{2}\cdot \cfrac{1}{2}}\)

OpenStudy (anonymous):

\[\sqrt{2-\sqrt{2}/}4 \]

OpenStudy (jdoe0001):

yeap, can't simplify it further